# Forecasting the future development in quality and value of professional football players

### v2

@@ 3.3 Feature importances
To illustrate the advantages of using explainable models, the feature importances are studied in both predictive problems. Among the models in this paper, the linear and tree-based models have methods that can be used to calculate the feature importances of the models. To investigate what features are important for the development of professional players, the feature importances are calculated. This is similar to the methods by Baouan et al. [47]. However, because the development of the indicator is considered instead of the indicator value itself, it is possible to describe the influence of the indicator value on the development. Min-max scaling is applied to the feature importances to be able to compare them across the different models.
@@ 4.1.1 General population of players
The loss values on the test set are shown in Figure 5 and Figure 5. The RMSE and MAE agree on the predictive performance of the models to forecast player development in general player quality, as they show similar patterns. The results indicate that the XGBoost model attains the lowest loss values, with the random forest having the second-best losses.
Figure 4: The RMSE loss per model in the player quality prediction problem.
Figure 5: The MAE loss per model in the player quality prediction problem.
@@ 4.1.2 Predictions per age
The RMSE loss is also determined for the players of each age in the test set as visualized in Figure 6. The XGBoost and random forest models perform better for all ages. For ages above 25, the XGBoost model has better loss values compared to the random forest model, whereas the differences are less evident for the lower ages. Because the younger ages are most important for estimating the potential of players in practice, these results indicate that both the XGBoost and random forest models are favorable models.
In general, Figure 6 shows that all models predict best on players with ages 24 up to 28. The quality of younger or older players is more volatile because young players tend to increase in quality, and older players tend to decrease in quality. Additionally, more data is available on players around their peak age. The worse predictive performance on the younger and older players is explained by the lower number of data points combined with a higher volatility in player quality.
Additionally, the results show that the kNN models perform worse on young players. The linear methods tend to perform worse for older players. This indicates that nonlocal methods with more features beyond time series perform better for young ages, whereas the patterns for older players involve nonlinearities or interaction effects. This explains the low loss values of the tree-based models on the general test set as shown in Figure 5 and Figure 5.
@@ 4.1.3 Prediction on important player groups
The test losses are determined on three different groups of players that are important for the application of the models in this prediction problem: high-quality players, players with a large decrease in performance, and players with a large improvement in performance. The RMSE values on the test set for these subgroups are shown in Figure 7, Figure 9, and Figure 9. In general, the vertical scales show that the RMSE on these subgroups is larger than that of the general population. This is even more evident for players with a large increase, and it indicates that the development of these players of interest is harder to predict, especially the players with a large increase.
The results also show that the tree-based models are best at predicting the development of all three of these subgroups of players. The random forest and XGBoost models appear to predict the best for the high-quality players and the players with a large decrease in quality. Figure 9 shows that the XGBoost model outperforms all other models on the group of players with a large increase in quality. The random forest model attains the second-best loss values.
Figure 7: The RMSE loss per model for players with a SciSkill of at least 100 in the player quality prediction problem.
Figure 8: The RMSE loss per model for players with a decrease of at least 10 in the player quality prediction problem.
Figure 9: The RMSE loss per model for players with a large increase of at least 10 in the player quality prediction problem.
@@ 4.1.4 Feature importances
The feature importances of the linear and tree-based models are shown in Figure 10. The feature importances indicate that the age (‘age_years’), age squared (‘age_years_squared’), and the difference between the age and the peak age (‘years_diff_peak_age’) are important features amongst the models. It is known that young players tend to improve in quality, whereas older players often decrease in quality. Therefore, it is reasonable that these age-related factors are indeed important in the development of the quality of professional football players.
Figure 10: The feature importances of the linear and tree-based models in the prediction problem of player quality. Min-max scaling has been applied to the feature importances.
Next to that, other features have also been assigned a large importance by the models. For instance, the player quality at prediction time (‘sciskill’) has a large feature importance. This indicates that the development patterns depend on the player level, which is reflected by the larger loss values in Figure 7 compared to Figure 5. The importance of the difference between player quality and average team quality (‘sciskill_diff_mean_team’) can be explained by the fact that players tend to have good performances when their team plays well. Consequently, the player’s quality grows towards the average team quality, which explains the importance of the difference in quality between the player and his team. Lastly, the feature importance assigns a large importance to the number of months since the last registered game (‘previous_zero_months’). The SciSkill model penalizes players when they have not played games for a long time. As this penalty is applied after the next game of a player, the quality of a player can be expected to decrease when the number of months since the last game is large. The importance of the number of months since the last registered game shows that the models can find this pattern.
@@ 4.1.5 Predictive accuracy
The results of training the models to predict the future player quality show that the XGBoost model provides the most accurate predictions on the general population of players. The XGBoost model outperforms the other models on young players, high-quality players, and players with either a large decrease or increase in football quality. The random forest model predicts second best, with similar performances on the young players and most of these interesting subgroups of players. For the problem of predicting a football player’s development in quality, it is concluded that the best predictive performance is attained by the XGBoost model, followed closely by the random forest model.
@@ 4.2.1 General population of players
The loss values for predicting player value development in the prediction problem of the monetary player value are given in Figure 12 and Figure 12. Although the differences in RMSE are less obvious than the differences in the MAE, the results show that the random forest model attains the lowest loss values for both the RMSE and MAE. The XGBoost model has the second-best values.
Figure 11: The RMSE loss per model in the player value prediction problem.
Figure 12: The MAE loss per model in the player value prediction problem.
@@
Additionally, the kNN models, which have time-series information as features, predict better than the linear models. These results show that the time series of the player performance indicators contains important information for the development of the transfer value and that nonlinearities and interactions are involved.
@@ 4.2.2 Predictions per age
The RMSE values on the test set for each age are given in Figure 13. The results show that the loss values of the models are similar at all ages and that there is no model that outperforms the other models on players of a young age. Therefore, no model is favorable based on the prediction of young players.
@@
Moreover, the linear models perform worse on older players. This suggests that the patterns in the development of older players involve interaction effects or nonlinearities that were not captured by the linear models.
Generally, the models tend to predict the future development of player value worse for young players and better for older players. The development of a young player is less predictable, similarly as found for the prediction of player quality in subsection 4.1. This is due to the fact that the KPIs for young players are more volatile. On the other hand, the player values of older players are smaller and more predictable because old players tend to decrease in value. In this way, it can be explained why the models have large loss values for young players and small loss values for older players.
@@ 4.2.3 Prediction on important player groups
For the development of the transfer value, the four different important groups of players that have been identified as important for the application of the models are: high-quality players, high-value players, players with a large decrease in transfer value, and players with a large improvement in transfer value. The RMSE for these groups of players is visualized in Figure 15, Figure 15, Figure 17, and Figure 17. Similar to the prediction problem for player quality, the errors in the predictions are larger for these interesting groups of players.
Figure 14: The RMSE loss per model for players with a SciSkill of at least 100 in the player value prediction problem.
Figure 15: The RMSE loss per model for players with a player value (ETV) of at least €10,000,000 in the player value prediction problem.
Figure 16: The RMSE loss per model for players with a decrease of at least €2,500,000 in the player value prediction problem.
Figure 17: The RMSE loss per model for players with an increase of at least €2,500,000 in the player value prediction problem.
The results in Figure 15 and Figure 17 show that the random forest model and kNN-based models perform best for the players of high quality and the players with a large decrease in value. The low loss values of the kNN-based models indicate that the time series features described in LABEL:Appendix:_Table_knn_features_player_value contain most of the important information to predict the development of these subsets of players. The good performance of the kNN also indicates that a local method is suitable for the prediction of these types of players. On the other hand, Figure 17 shows that, for players with a large increase, the kNN models perform relatively worse compared to the tree-based models. Similarly, the tree-based models also outperform the linear model on the subset of players with a large increase. These differences suggest that a large increase in player value is influenced by many different features and contains nonlinear patterns or interaction effects.
Figure 15 shows that the performance of the random forest model on the players with a high player value is evidently better than that of the other models. Players often attain high transfer values for a short period in their careers when they show peak performance and are young. Consequently, players with high transfer values can sometimes be expected to decrease in value shortly after obtaining high transfer values. The random forest model seems to be the only model that captures this pattern, as will be indicated by the high feature importance of the current transfer value in subsubsection 4.2.4. This explains the significantly better performance of the random forest model on the high-value players.
Generally, the random forest model attains relatively low loss values on the interesting subgroups of players. This makes the random forest favorable over the other models. Additionally, the kNN-based models perform well on high-quality players and players with a large decrease, making them more suitable to predict star players on their decline.
@@ 4.2.4 Feature importances
The feature importances for the development of the player transfer value after min-max scaling are shown in Figure 18. The results indicate that the most important features for predicting player value development are the features describing the most recent transfer value and the developments within the last 6 and 12 months of the player’s transfer value. Especially, the random forest model assigns a large value to the player value at the time of prediction (‘etv’), which explains its good performance on high-value players in Figure 15. Figure 18 also shows that additional time series information about the player quality (‘sciskill_diff_6m_ago’) provides information about the development of the player transfer value. This implies that the time series information is most important in predicting the future development, which is in line with the relatively good predictive performance of the kNN models that are based on time series information.
Figure 18: The feature importances of the linear and tree-based models in the prediction problem of player value. Min-max scaling has been applied to the feature importances.
Moreover, it is found that the month of the year is an important factor. This can be explained by the fact that the transfer value of a player develops differently at different times of the year. For example, a player’s transfer value tends to decrease when the player only has 6 months of contract left because he could negotiate with other professional football clubs. As this commonly happens in the winter and the Estimated Transfer Value model contains the indicator whether a player has less than 6 months on his contract, the development of players’ transfer value is different depending on the time of the year.
@@ 4.3 Predictive accuracy
The results showed that the random forest model attained the best loss values on the general population of players, the groups of players with high quality, high value, or with a large increase or decrease in player value. To conclude, it is found that the random forest model provides the most accurate predictions of the development of a football player’s monetary value.
5 Conclusion
This paper aims to find the most suitable supervised learning model for forecasting the development of player performance indicators one year ahead. Through the literature, it was found that linear regression, random forest, and kNN-based models are favorable based on their uncertainty quantification methods. Two prediction problems have been considered to study the predictive performance of linear, tree-based, and kNN-based models.
The results show that the predictive performance of the random forest has the second-best predictive accuracy for predicting the development of the player quality and that it provides the best predictive accuracy for the forecasting of the player value. Moreover, an off-the-shelf method for uncertainty quantification is available. Therefore, it is concluded that the random forest is the most suitable explainable machine learning model to predict the development of player performance indicators one year ahead.
@@ 6 Implications and future directions
By addressing the two prediction problems, two random forest models have been obtained that can predict the development of both the quality and the monetary value of football players. These random forest models can be used to aid in transfer decisions. Suppose the manager of a football club has a long-term interest in a football player. If the model predicting the development of player quality indicates that a player will grow in quality, it means that a player is more interesting to a football club in the long term. If the prediction of the transfer value indicates an increase in the transfer value, it might be better to buy the player sooner rather than later. The two models can also be combined to give extra insights. Suppose a manager currently has a veteran player who is predicted to start decreasing in quality, and he can buy a young player who is predicted to develop into a first-team player within the next year. Assume that the transfer value of the young player is predicted to only slightly increase. In this case, it might be better to buy the young player one year later and then sell the veteran player, because the young player will only have a slightly higher transfer fee, and the veteran player will have a better quality in the meantime. These examples illustrate the added value of our models to predict player development for the improvement of data-informed decision-making.
@@
The models from this paper can also be used to complement existing methods in the literature. Pantuso and Hvattum [19] introduced a method to optimize transfer decisions based on indicators that describe a player’s quality and transfer value and their future values. Their methodology needs to know the ‘future’ values of the player quality and their transfer values. To solve this, they consider transfer situations from over a year prior so that the values of one year later are already known. Consequently, their model can only be applied to historical situations. By using our models to predict the future values in quality and transfer value, it is possible to obtain these predictions and optimize transfer decisions with a current-time application. This makes it possible to advise data-driven transfer decisions to optimize the squad in a real-life transfer period.
@@
An advantage of the explainable models in this research is that they provide methods to gain insight into important factors for prediction. The models for the prediction problems show that this can be done via feature importance, similar as done by Baouan et al. [47]. Because our models predict the difference between the current performance indicators and those of one year later, our research can also show the influences of the indicator itself. Our results show that the value of the indicator and the historical values of the indicators are the most important features, which adds to the knowledge obtained by Baouan et al. [47] that describes which features are important. Additionally, time-dependent variables like the period in the year and the months without games have been found to be important features in forecasting player development. This indicates that the time series of the indicators themselves contains important information on the development of football players, although our findings also suggest that contextual information gives improved predictive performance of the models.
@@
Acknowledgements The authors thank Geurt Jongbloed for his comments on a draft version of this paper and the company SciSports for the data, computational resources, and insight into the needs of practitioners.
@@
Author contributions K.A. carried out the research with supervision and input from F.G. and J.S.

### v3

@@ 3.3 Feature Importances
To illustrate the advantages of using explainable models, the feature importances are studied in both predictive problems. Among the models in this paper, the linear and tree-based models have methods that can be used to calculate the feature importances of the models. To investigate what features are important for the development of professional players, the feature importances are calculated. This is similar to the methods by Baouan et al. (2022). However, because the development of the indicator is considered instead of the indicator value itself, it is possible to describe the influence of the indicator value on the development. Min–max scaling is applied to the feature importances to be able to compare them across the different models.
@@ 4.1.1 General Population of Players
The loss values on the test set are shown in Figure 4. The RMSE and MAE agree on the predictive performance of the models to forecast player development in general player quality, as they show similar patterns. The results indicate that the XGBoost model attains the lowest loss values of all models. The decision tree and random forest models also attained low loss values, with the latter having slightly better loss values.
@@
Figure 4: The test loss values for the different models on the general population of players in the player quality prediction task for two different loss functions: (a) RMSE and (b) MAE.
@@ 4.1.2 Predictions per Age
The RMSE loss is also determined for the players of each age in the test set, as visualized in Figure 5. In general, Figure 5 shows that all models predict best on players with ages 24 up to 28. The quality of younger or older players is more volatile because young players tend to increase in quality, and older players tend to decrease in quality, albeit over a longer time period. Additionally, more data is available on players around their peak age. The data set consists of roughly 10,000 data points for players aged 18 and 20,000 for players aged 34, while around 70,000 data points describe situations of players aged 25. The worse predictive performance on the younger and older players can be explained by the lower number of data points combined with a higher volatility in player quality. Moreover, the low number of data points for young players indicates that the RMSE estimates for young players may be less stable than those of players aged 24 up to 28.
@@
The results show that the XGBoost model attains the lowest loss values for almost all ages, especially for ages above 22. For the ages below 22, the random forest model attains similar loss values. Because the younger ages are most important for estimating the potential of players in practice, these results indicate that both the XGBoost and random forest models are favorable models.
Additionally, the results show that the kNN models perform worse on young players. Generally, younger players are harder to model because of their volatility. Combined with the fewer available data points, this explains why the local kNN models can predict the development of younger players less well. Conversely, the linear methods tend to perform worse for older players. This indicates that the patterns for older players involve nonlinearities or interaction effects. Because tree-based models are non-local methods that can take into account nonlinearities and interaction effects, they do not suffer from these problems. This explains the low loss values of the tree-based models on the general test set as shown in Figure 4a,b.
@@ 4.1.3 Prediction on Important Player Groups
The test losses are determined on three different groups of players that are important for the application of the models in this prediction problem: high-quality players, players with a large decrease in performance, and players with a large improvement in performance. Based on domain knowledge, high-quality players were defined as players with a SciSkill of more than 100, while an increase or decrease of more than 10 was defined as large. The RMSE values on the test set for these subgroups are shown in Figure 6. In general, the vertical scales show that the RMSE on these subgroups is larger than that of the general population. This is even more evident for players with a large increase, and it indicates that the development of these players of interest is harder to predict.
The results also show that the tree-based models are best at predicting the development of all three of these subgroups of players. The XGBoost model attains the best RMSE values for all three of the important groups of players. Figure 6c shows that the XGBoost model outperforms all other models on the group of players with a large increase in quality, whereas the differences in loss values with the decision tree and the random forest are less apparent in Figure 6a,b.
@@ 4.1.4 Feature Importances
The feature importances of the linear and tree-based models are shown in Figure 7. The feature importances indicate that the age (‘age_years’), age squared (‘age_years_squared’), and the difference between the age and the peak age (‘years_diff_peak_age’) are important features amongst the models. These features are strongly correlated, as they all capture age-related effects. As a consequence, the models often identify at most two of these features as important. The importance of age-related features is in line with domain knowledges since it is known that young players tend to improve in quality, whereas older players often decrease in quality. Therefore, it is reasonable that these age-related factors are indeed important in the development of the quality of professional football players.
Figure 6: The RMSE loss per model in the player quality prediction task for various player subsets: (a) players with a SciSkill of at least 100, (b) players with a SciSkill decrease of at least 10, and (c) players with a SciSkill increase of at least 10.
Next to that, other features have also been assigned a large importance by the models. For instance, the player quality at prediction time (‘sciskill’) has a large feature importance. This indicates that the development patterns depend on the player level, which is reflected by the larger loss values in Figure 6a compared to Figure 4a. The importance of the difference between player quality and average team quality (‘sciskill_diff_mean_team’) can be explained by the fact that players tend to have good performances when their team plays well. Consequently, the player’s quality grows towards the average team quality, which explains the importance of the difference in quality between the player and his team. Lastly, the feature importance assigns a large importance to the number of months since the last registered game (‘previous_zero_months’). The SciSkill model penalizes players when they have not played games for a long time. As this penalty is applied after the next game of a player, the quality of a player can be expected to decrease when the number of months since the last game is large. The importance of the number of months since the last registered game shows that the models can find this pattern.
Figure 7: The feature importances of the linear and tree-based models in the prediction problem of player quality. Min–max scaling has been applied to the feature importances.
@@ 4.1.5 Predictive Accuracy
The results show that the tree-based models tend to attain the lowest loss values when predicting the development of a player’s quality in this specific prediction problem. The XGBoost model seems to provide the most accurate predictions for the general population of players in the data set. Moreover, the XGBoost model attains low loss values on young players, high-quality players, and players with either a large decrease or increase in football quality in this data set. The random forest model attains similar performances on the young players and most of the interesting subgroups of players. Overall, the XGBoost model seems to have the lowest loss values on the data set for predicting player quality, followed by the random forest model.
@@ 4.2.1 General Population of Players
The loss values for predicting player value development in the prediction problem of the monetary player value are given in Figure 8. The results show that the tree-based and kNN-based models perform relatively well in this prediction problem. Although the differences in RMSE are less obvious than the differences in the MAE, the results show that the random forest model attains the lowest loss values for both the RMSE and MAE, while the XGBoost and kNN-based models attain only slightly higher loss values.
Figure 8: The test loss values for the different models on the general population of players in the player value prediction task for two different loss functions: (a) RMSE and (b) MAE.
@@
Additionally, the kNN models, which have time series information as features, predict better than the linear models. These results show that the time series of the player performance indicators contains important information for the development of the transfer value and that nonlinearities and interactions are involved.
@@ 4.2.2 Predictions per Age
The RMSE values on the test set for each age are given in Figure 9. The results show that the loss values of the models are similar at all ages and no clear distinction can be made between the quality of the models based on this.
@@
Moreover, the linear models perform worse on older players. This suggests that the patterns in the development of old players involve interaction effects or nonlinearities that were not captured by the linear models.
Generally, the models tend to predict the future development of player value worse for young players and better for older players. The development of a young player is less predictable, similarly as found for the prediction of player quality in Section 4.1. This is due to the fact that the KPIs for young players are more volatile. On the other hand, the player values of older players are smaller and more predictable because older players tend to decrease in value. In this way, it can be explained why the models have large loss values for young players and small loss values for older players.
@@ 4.2.3 Prediction on Important Player Groups
For the development of the transfer value, the four different groups of players that have been identified as important for the application of the models are high-quality players, high-value players, players with a large decrease in transfer value, and players with a large improvement in transfer value. Based on domain knowledge, high-quality players were defined as players with a SciSkill of more than 100, while high-value players are defined as players with an ETV of at least 10 M EUR. An increase or decrease of more than EUR 2.5 M was defined as large. The RMSE for these groups of players is visualized in Figure 10. Similar to the prediction problem for player quality, the errors in the predictions are larger for these interesting groups of players.
Figure 10: The RMSE loss per model in the player value prediction task for various player subsets: (a) players with a SciSkill of at least 100, (b) players with an ETV of at least EUR 10M, (c) players with a value decrease of at least EUR 2.5M, and (d) players with a value increase of at least EUR 2.5M.
The results in Figure 10a,c show that the random forest model and kNN-based models attain the lowest loss values for the players of high quality and the players with a large decrease in value. The low loss values of the kNN-based models indicate that the time series features, described in Table S6 in the Supplementary Materials, contain most of the important information to predict the development of these subsets of players. The good performance of the kNN also indicates that a local method is suitable for the prediction of these types of players. On the other hand, Figure 10d shows that, for players with a large increase, the kNN models attain worse loss RMSE values compared to the tree-based models. Similarly, the tree-based models also outperform the linear model on the subset of players with a large increase. These differences suggest that a large increase in player value is influenced by many different features and contains nonlinear patterns or interaction effects.
Figure 10b shows that the random forest model has a distinctly lower RMSE on the players with a high player value than the other models. Players often attain high transfer values for a short period in their careers when they show peak performance and are young. Consequently, players with high transfer values can sometimes be expected to decrease in value shortly after obtaining high transfer values. The random forest model seems to be the only model that captures this pattern, as will be indicated by the high feature importance of the current transfer value in Section 4.2.4. This explains the lower loss values of the random forest model on the high-value players.
@@ 4.2.4 Feature Importances
The feature importances for the development of the player transfer value after min–max scaling are shown in Figure 11. The results indicate that the most important features for predicting player value development are the features describing the most recent transfer value and the developments within the last 6 and 12 months of the player’s transfer value. Especially the random forest model assigns a large value to the player value at the time of prediction (‘etv’), which explains its good performance on high-value players in Figure 10b. Figure 11 also shows that additional time series information about the player quality (‘sciskill_diff_6m_ago’) provides information about the development of the player transfer value. This implies that the time series information is most important in predicting the future development, which is in line with the relatively good predictive performance of the kNN models that are based on time series information.
Moreover, it is found that the month of the year, which indicates whether it is January or July, is an important factor. Transfers in the winter transfer period are often driven by a necessity of a specific type of player at that moment, while transfers to improve the squad in the long term are more frequent in the summer transfer period. This changes the transfer fees, which is reflected in the ETV values. Additionally, players with only 6 months of contract left are able to negotiate a free transfer with a club, which decreases the transfer values of these players. As this commonly happens in the winter and the Estimated Transfer Value model contains the indicator whether a player has less than 6 months on his contract, the development of players’ transfer value is different depending on the time of the year. Therefore, the fact that the month of the year is an important factor is in line with domain knowledge.
@@ 4.3 Predictive Accuracy
The results showed that, with this feature set, the random forest, XGBoost, and kNN-based models attained low loss values on the general population of football players. No clear distinction was found between the RMSE values for players of different ages. The random forest and kNN models attained low RMSE scores on the subsets of high-quality players and players with a large decrease in monetary value. The tree-based models had the lowest loss values on players with a large increase in value, with the XGBoost model attaining the lowest loss. The random forest model attained a lower loss value on the subset of high-value players. Taking this into account, the random forest model has the best predictive accuracy on the data set of this prediction problem, while the XGBoost model and the kNN-based models also attain low loss values.
Figure 11: The feature importances of the linear and tree-based models in the prediction problem of player value. Min–max scaling has been applied to the feature importances.
5 Conclusions
This paper aims to find the most suitable supervised learning model for forecasting the development of player performance indicators one year ahead in a practitioner-oriented way. Through the literature, it was found that linear regression, random forest, and kNN-based models are favorable based on their uncertainty quantification methods.
Two prediction problems were considered to study the predictive performance of linear, tree-based, and kNN-based models. The results show that the XGBoost model attained the lowest loss values for predicting the development of the player quality with this data set, while the random forest also had relatively low loss values. For the prediction of the player value, it was found that the random forest had the lowest loss values. The XGBoost and kNN-based models also had relatively accurate predictions according to some of the studied losses.
Because the random forest had a good predictive accuracy on these data sets and it provides methods for uncertainty quantification, it seems to be a suitable model for predicting the development of player quality and value in football. Nonetheless, it is important to note that random forests can be sensitive to data imbalance, which regularly occurs in elite football. Additionally, given the large number of features needed to predict the development of football players, random forests may risk overfitting if not properly tuned and validated. These limitations should be considered when applying the model in practice, and appropriate techniques, like resampling or feature selection, may be necessary to mitigate them.
Taking these considerations into account, it is concluded that the random forest is the most suitable explainable machine learning model to predict the development of player performance indicators one year ahead for the above prediction problems.
@@ 6 Implications and Future Directions
By addressing the two prediction problems, two random forest models have been obtained that can predict the development of both the quality and the monetary value of football players. These random forest models can be used to aid in transfer decisions combined with a critical view of a domain expert. Suppose the manager of a football club has a long-term interest in a football player. If the model predicting the development of player quality indicates that a player will grow in quality, it means that a player is more interesting to a football club in the long term. If the prediction of the transfer value indicates an increase in the transfer value, it might be better to buy the player sooner rather than later. The two models can also be combined to give extra insights. Suppose a manager currently has a veteran player who is predicted to start decreasing in quality, and he can buy a young player who is predicted to develop into a first-team player within the next year. Assume that the transfer value of the young player is predicted to only slightly increase. In this case, it might be better to buy the young player one year later and then sell the veteran player after a critical evaluation of the possible influences of noise and expected transfer market development by the manager. This would be beneficial because the young player will only have a slightly higher transfer fee, and the veteran player will have a better quality in the meantime. These examples illustrate the added value of our models to predict player development for the improvement of data-informed decision-making.
@@
The models from this paper can also be used to complement existing methods in the literature. Pantuso and Hvattum (2021) introduced a method to optimize transfer decisions based on indicators that describe a player’s quality and transfer value and their future values. Their methodology needs to know the ‘future’ values of the player quality and their transfer values. To solve this, they consider transfer situations from over a year prior so that the values of one year later are already known. Consequently, their model can only be applied to historical situations. By using our models to predict the future values in quality and transfer value, it is possible to obtain these predictions and optimize transfer decisions with a current-time application. This makes it possible to advise data-driven transfer decisions to optimize the squad in a real-life transfer period.
When applying the models from this research at football clubs, possible shortcomings should be taken into consideration. The models will most likely capture patterns of development for players that are typical for the population on which they are trained. They will perform less on nontypical players, which makes it harder to predict, e.g., late bloomers. Young or injury-prone players can be considered vulnerable players who are nontypical for the general population, and the predictions of the models on such players should thus be interpreted with extra care. This highlights the value of explainable modeling, which can be complemented with interpretation techniques like the use of Shapley values by Lundberg and Lee (2017).
@@
Another advantage of the explainable models in this research is that they provide methods to gain insight into important factors for prediction. The models for the prediction problems show that this can be achieved via feature importance, similar as carried out by Baouan et al. (2022). Because our models predict the difference between the current performance indicators and those of one year later, our research can also show the influences of the indicator itself. Our results show that the value of the indicator and the historical values of the indicators are the most important features, which adds to the knowledge obtained by Baouan et al. (2022) that describes which features are important. Additionally, time-dependent variables like the period in the year and the months without games have been found to be important features in forecasting player development. This indicates that the time series of the indicators themselves contains important information on the development of football players, although our findings also suggest that contextual information gives improved predictive performance of the models.